Aujuanaih235:

Maria rewrites a fraction less than 1 as a decimal. The numerator is a whole number greater than 0. For which denominator will the fraction always convert to a repeating decimal? 15 18 6

KIllerwolfdaone:

think about what a repeating decimal is and what ever is the best answer you can pick it and recheck to make sure

surjithayer:

actually all can give repeating decimal. 2/15=0.1333333... 5/18=0.277777... 1/6=0.1666666...

Vocaloid:

There was a missing answer choice that the OP didn't post (it was a prime number, which is the only type of integer that will guarantee a repeating decimal because it won't divide evenly for any integers except itself, and 1, neither of which produce a fraction < 1)